# Language Models are Injective and Hence Invertible

### v3

@@ A.1 Notation
Matrices (vectors) are denoted by uppercase (lowercase) bold-face letters: $\mathbf{X}\in\mathbb{R}^{d_{1}\times d_{2}}$ ($\mathbf{x}\in\mathbb{R}^{d}$). For vectors and matrices, we frequently use standard norms and common matrix operations. The Hadamard and Kronecker products are defined following Kolda & Bader (2009):
@@ A.2.2 Real-Analytic Functions with Matrix Inputs
Let $\mathrm{vec}_{m,n}:\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$ denote the standard vectorization operator, which stacks the columns of a matrix into a single column vector (Henderson & Searle, 1981).
We also define the corresponding matricization operator $\mathrm{mat}_{m,n}:\mathbb{R}^{mn}\to\mathbb{R}^{m\times n}$. As shown in Chacón & Duong 2020, the vectorization and matricization operators are mutual inverses:
@@ A.2.3 Real Analyticity of Common Components
We now collect several building blocks that will be used repeatedly. Throughout, all maps are defined on $\mathbb{R}^{m\times n}$, an open set, so Definition A.2 applies.
@@
The exponential map $\exp:\mathbb{R}\to(0,\infty)$ is real-analytic with $\exp^{\prime}(y)\neq 0$ for all $y$. By the real-analytic inverse function theorem (see Krantz & Parks 2002, Thm. 2.3.1), its local inverse $\log$ is real-analytic on $(0,\infty)$. ∎
@@ A.3 Differential, Measure-Theoretic, and Topological Tools
Definition A.6 (Second Fréchet derivative (Magnus & Neudecker, 2019, Ch. 18)).
@@
as noted in Magnus & Neudecker 2019, Ch. 18.
@@ C.2 Absolute continuity of the parameter distribution is preserved under GD
Our goal in this subsection is to explain why absolute continuity of the parameter law at initialization survives any finite number of gradient–descent (GD) steps, thereby allowing the almost-sure injectivity argument from the previous subsection to persist throughout training. The story begins with regularity: by Proposition B.3 and Proposition A.6, the loss $\mathcal{L}_{\mathrm{s},\mathbf{p}}$ is real-analytic, and real-analyticity is closed under differentiation and composition. Consequently the GD map $\phi(\bm{\theta})=\bm{\theta}-\eta\nabla\mathcal{L}_{\mathrm{s},\mathbf{p}}(\bm{\theta})$ is real-analytic, its Jacobian $D\phi(\bm{\theta})=\mathbf{I}_{p}-\eta\nabla^{2}\mathcal{L}_{\mathrm{s},\mathbf{p}}(\bm{\theta})$ is real-analytic, and so is $\bm{\theta}\mapsto\det D\phi(\bm{\theta})$ (the determinant is a polynomial in the matrix entries). We then rule out the degenerate case by a witness: at $\bm{\theta}^{\star}=\mathbf{0}_{p}$, our Hessian calculation (Lemma C.4) shows $\det D\phi(\bm{\theta}^{\star})>0$, hence $\det D\phi$ is not identically zero and its zero set $\mathcal{C}:=\{\det D\phi=0\}$ has Lebesgue measure zero by the real-analytic zero–set theorem (Theorem A.1; summarized in Theorem C.3). On the complement $\mathbb{R}^{p}\setminus\mathcal{C}$, the Inverse Function Theorem (Theorem A.2) provides, for every $\bm{\theta}$, a neighborhood on which $\phi$ is a $C^{1}$ diffeomorphism. Although these neighborhoods form an a priori uncountable cover, the second countability of $\mathbb{R}^{p}$ (and of its subspaces) ensures a countable subcover of such charts (Proposition A.15, Lemma C.5). This countability is crucial because it lets us pass from local statements to a global measure statement via countable unions. With this cover in hand, the change-of-variables formula on each chart (Theorem C.4) implies that the image under the local inverse of any null set remains null; piecing the charts together and adding the null set $\mathcal{C}$ shows that preimages of Lebesgue-null sets under $\phi$ are null (Lemma C.6). Equivalently, $\phi$ pushes absolutely continuous laws to absolutely continuous laws (Theorem C.5); iterating across finitely many GD steps preserves absolute continuity (Corollary C.5.1). Finally, combining this preservation with the almost-sure pairwise distinctness of last-token representations over any finite input family (Corollary C.2.1) yields the main consequence we need for training: the last-token representation map remains injective almost surely after any finite GD horizon.
@@ C.2.1 Witness Construction
In particular, $Df(\mathbf{v}_{0})\big[(\mathbf{0}_{m},\;\cdot\;)\big]=\mathbf{0}_{n}$. The second-order chain rule for Fréchet derivatives (e.g. Magnus & Neudecker 2019, Thm. 18.4) yields:
@@
The determinant of a block triangular (in particular block diagonal) matrix equals the product of the determinants of its diagonal blocks (e.g. Horn & Johnson 2013, Cor. 0.8.5). Hence
@@
Lemma C.3 (Hessian of $\mathcal{L}$ w.r.t. $\mathbf{U},\bm{\beta}$ at $\bm{\theta}^{\star}=\mathbf{0}$ and its spectrum).
@@
Using $\mathrm{vec}(\mathbf{A}\mathbf{X}\mathbf{b})=(\mathbf{b}^{\top}\otimes\mathbf{A})\,\mathrm{vec}(\mathbf{X})$ (standard identity for vectorization, cf. Henderson & Searle (1981)), with $\mathbf{A}=\mathbf{I}_{n}$ and $\mathbf{b}=\bm{\beta}$,
@@
Combining, Lemma C.2 and Lemma C.3, we get that
$\mathrm{spec}\big(\nabla^{2}\mathcal{L}(\bm{\theta}^{\star})\big)$ $=\mathrm{spec}\big(\nabla^{2}_{(\mathbf{u},\bm{\beta})}\mathcal{L}(\bm{\theta}_{\star})\big)\ \cup\ \{0\}^{\,d+p^{\prime}}$
@@
with stepsize $\eta\in(0,1)$. Then the critical set
@@
By Proposition B.3, Proposition A.6 and the closure properties of real analyticity, $\mathcal{L}_{\mathrm{s},\mathbf{p}}$ is real-analytic; hence so are its gradient and Hessian. Therefore $\phi$ is real-analytic (Lewis, 2014, Thm. 1.1.15) and
@@
$\det D\phi(\bm{\theta}^{\star})=\left(1-\eta^{2}\|\mathbf{w}\|_{2}^{2}\right)^{d}>0.$
@@ C.2.2 Gradient Descent preserves absolute continuity
1) $\bm{\mathcal{X}}$ is open: By Proposition B.3, Proposition A.6 and the closure rules of real-analyticity, $\mathcal{L}_{\mathrm{s},\mathbf{p}}$ is $C^{2}$, hence $\phi$ is $C^{1}$. The map $\bm{\theta}\mapsto D\phi(\bm{\theta})$ is continuous, and the determinant is a continuous polynomial in the entries, so $g(\bm{\theta}):=\det D\phi(\bm{\theta})$ is continuous. Therefore $\mathcal{C}=g^{-1}(\{0\})$ is closed (Rudin, 1976, Thm. 4.8) and $\bm{\mathcal{X}}=\mathbb{R}^{p}\setminus\mathcal{C}$ is open.
@@
Then, gradient-descent preserves absolute continuity: for every absolutely continuous probability law $\mu$ on $\mathbb{R}^{p}$, its image under $\phi$ remains absolutely continuous:
@@
Appendix D Left-Invertibility Via SIP-It
Goal. We study when and how the hidden states of a causal decoder-only Transformer admit a left inverse: given the layer-$\ell$ representation at position $t$ and the true prefix $\pi=\mathrm{s}_{1:t-1}$, can we recover the next token $\mathrm{s}_{t}$?
Main idea. Under mild randomness in the parameters and causal masking, the one-step last-token map that sends a candidate token $v$ to the layer-$\ell$ representation at position $t$ (conditioning on $\pi$) is almost-surely injective, and in fact has a positive separation margin. This yields a simple verifier: declare $v$ correct iff the observed hidden state lies in a small ball around $F(v;\pi,t)$.
Algorithmic consequence. Because causality localizes the dependence to $(\pi,\mathrm{s}_{t})$, we can invert an entire sequence sequentially with a single pass over the vocabulary per position. We call this procedure Sip-It (Sequential Inversion via Prefixwise Injective Tests), and we show exact (and robust) recovery holds almost surely, with worst-case time $\Theta(T|\mathcal{V}|)$.
@@ Appendix D Left-Invertibility Via SIP-It
Every attention layer in every block is causal in the sense of Definitions B.6/B.7. Consequently, for any $\mathrm{s}$ and any $t\in[T]$,
@@
Sip-It is applied to models initialized with parameters drawn from an absolutely continuous distribution and trained via (mini-batch) gradient descent with step sizes in $(0,1)$, as described in Appendix C. Under these conditions, any network considered in the sequel is almost-surely injective (Theorem C.1).
@@ D.2 The Core Routines: Local Verifiers, Acceptance Regions, and Policies
Finally, we introduce the last conceptual block required to build the inversion algorithm:
@@
Both Alg. 2 and Alg. 3 satisfy Definition D.3. In practice we use the gradient-guided policy with standard gradient descent updates, as it tends to find the verified token with far fewer proposals: the next token is chosen by ranking $\mathcal{V}$ by the distance $\|\mathbf{E}_{v}-\mathbf{e}^{(j)}\|_{2}$ to the updated proxy $\mathbf{e}^{(j)}$. This preserves the same worst-case guarantees (single pass over $\mathcal{V}$) while improving empirical efficiency.
D.3 Global Inversion via Sip-It
@@
We are now ready to introduce our inversion algorithm: Sip-It (Alg. 1). The algorithms applies to decoder-only transformers with causal self-attention (Assumption D.1), and assumes injectivity, which occurs with almost-surely (Assumption D.2). We assume access to the layer-$\ell$ hidden states per position $\left\{\widehat{\mathbf{h}}_{t}\right\}_{t=1}^{T}$ and to the parameters needed to evaluate the local verifier from Definition D.2 for arbitrary $(t,\pi,j)$, as well as the gradient (when needed), namely to the model up to layer $\ell$. A policy algorithm is fixed (e.g., Alg. 3).
@@
Fix $t$ and a true prefix $\pi=\langle\mathrm{s}_{1},\ldots,\mathrm{s}_{t-1}\rangle$. Under Assumption D.1, the observed hidden state at step $t$ satisfies $\mathbf{h}_{t}(\mathrm{s})=F(\mathrm{s}_{t}\,;\,\pi,t)$ (Lemma D.2). In addition, under Assumption D.2, $F$ is injective and has positive margin $\Delta_{\pi,t}>0$ almost-surely (Theorem D.1 and Lemma D.1). Consequently, for the local verifier of Definition D.2, the following hold with probability one:
@@
Work on the probability-one event of Proposition D.2 (under Assumption D.1 and Assumption D.2 with the stated $\varepsilon$), on which the local verifier at step $t$ uniquely accepts the true token $\mathrm{s}_{t}$. Equivalently,
@@
Theorem D.2 (Correctness of Sip-It (noiseless & robust)).
For each $t\in\{1,\ldots,T\}$ let $\pi_{t}=\langle\mathrm{s}_{1},\ldots,\mathrm{s}_{t-1}\rangle$ and let $\Delta_{\pi_{t},t}>0$ be the margin of the one-step map $F(\cdot;\pi_{t},t)$ from Lemma D.1. Under Assumptions D.1 and D.2, run Sip-It (Alg. 1) with a tolerance $\varepsilon\geq 0$ and observations
@@
By Assumption D.2, Theorem D.1, and Lemma D.1, there is a probability-one event on which, for all $t$, $F(\cdot;\pi_{t},t)$ is injective with strictly positive margin $\Delta_{\pi_{t},t}$. Intersecting across finitely many $t$ preserves probability 1. Work on this event.
By Assumption D.1 and Lemma D.2, $\mathbf{h}_{t}(\mathrm{s})=F(\mathrm{s}_{t};\pi_{t},t)$. Since $\|\mathbf{e}_{t}\|_{2}\leq\varepsilon$,
@@
When $\varepsilon=0$, equation 36 also holds by Proposition D.1(1). We now analyze Sip-It and proceed by induction on $t$.
@@
Run Sip-It (Alg. 1) on a length-$T$ sequence with any policy that enumerates $\mathcal{V}$ without replacement. Then the algorithm halts after a finite number of iterations. Moreover, in the worst case the inner for-loop over $j$ executes at most $|\mathcal{V}|$ iterations at each position $t$, so the total number of verifier tests across the entire run is at most $T\,|\mathcal{V}|$. In particular, the number of loop iterations grows linearly with $T\cdot|\mathcal{V}|$.
@@
Appendix E Additional Experiments and Implementation Details

### v4

@@ A.1 Notation
Matrices (vectors) are denoted by uppercase (lowercase) bold-face letters: $\mathbf{X}\in\mathbb{R}^{d_{1}\times d_{2}}$ ($\mathbf{x}\in\mathbb{R}^{d}$). For vectors and matrices, we frequently use standard norms and common matrix operations. The Hadamard and Kronecker products are defined following Kolda and Bader (2009):
@@ A.2.1 Real-Analytic Functions with Vector Inputs
We begin with the standard vector-space definition and its basic algebraic properties. These are the building blocks from which all later analyticity arguments are assembled.
@@ A.2.2 Real-Analytic Functions with Matrix Inputs
Since transformer layers operate on matrices (e.g., $\mathbf{X}\in\mathbb{R}^{T\times d}$), we need to extend real-analyticity from vector spaces to matrix spaces. The key tool is the vectorization operator, which lets us reduce matrix-analytic questions to the vector case treated above.
@@
Let $\mathrm{vec}_{m,n}:\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$ denote the standard vectorization operator, which stacks the columns of a matrix into a single column vector (Henderson and Searle, 1981).
We also define the corresponding matricization operator $\mathrm{mat}_{m,n}:\mathbb{R}^{mn}\to\mathbb{R}^{m\times n}$. As shown in Chacón and Duong 2020, the vectorization and matricization operators are mutual inverses:
@@ A.2.3 Real Analyticity of Common Components
We now catalogue the specific functions that appear inside transformer layers, proving each one is real-analytic. These building blocks—polynomials, exponentials, softmax, row normalization, matrix products, Hadamard scaling, and stacking—will be composed in Appendix B to establish the real-analyticity of the full model. Throughout, all maps are defined on $\mathbb{R}^{m\times n}$ (or an open subset thereof), so Definition A.2 applies.
@@
The exponential map $\exp:\mathbb{R}\to(0,\infty)$ is real-analytic with $\exp^{\prime}(y)\neq 0$ for all $y$. By the real-analytic inverse function theorem (see Krantz and Parks 2002, Thm. 2.3.1), its local inverse $\log$ is real-analytic on $(0,\infty)$. ∎
The next three results handle the attention-specific operations: softmax, row normalization (used in the causal projection form), and entrywise matrix polynomials.
@@
Finally, we record the algebraic operations—matrix multiplication, Hadamard scaling, concatenation, and noncommutative matrix polynomials—that allow us to compose the above primitives into full transformer layers.
@@ A.3 Differential, Measure-Theoretic, and Topological Tools
Definition A.6 (Second Fréchet derivative (Magnus and Neudecker, 2019, Ch. 18)).
@@
as noted in Magnus and Neudecker 2019, Ch. 18.
Euclidean topology.
The following standard definitions and facts about $\mathbb{R}^{p}$ are used in the local-to-global measure argument in Appendix C.
@@
Invertibility and measure transport.
The inverse function theorem and the pushforward formalism are the two tools that connect local diffeomorphism charts to global measure-preservation statements.
@@ Appendix B Transformer Language Model
Input processing.
The first stage of the model maps a discrete token sequence into a continuous matrix representation via learned embeddings.
@@
Sub-layer modules.
The Transformer block is built from four reusable sub-layers—an MLP, (causal) self-attention, multi-head attention, and layer normalization—each defined next.
@@
Full architecture assembly.
With all sub-layers in place, we assemble them into a single pre-LN residual block, stack $L$ such blocks into the Transformer backbone, and append the unembedding head to form the complete language model.
@@
Verification of real-analyticity.
We close this section by showing that every module defined above is jointly real-analytic in its inputs and parameters. This is the technical property that lets the measure-zero arguments in Appendix C go through. We first record the equivalence between the two causal-softmax formulations, then verify analyticity of the embedding layer and of each sub-layer and their compositions.
@@ C.2 Absolute continuity of the parameter distribution is preserved under GD
Our goal in this subsection is to explain why absolute continuity of the parameter law at initialization survives any finite number of gradient–descent (GD) steps, thereby allowing the almost-sure injectivity argument from the previous subsection to persist throughout training. The argument proceeds in four steps.
Step 1: Regularity of the GD map.
By Propositions B.3 and A.6, the loss $\mathcal{L}_{\mathrm{s},\mathbf{p}}$ is real-analytic, and real-analyticity is closed under differentiation and composition. Consequently the GD map $\phi(\bm{\theta})=\bm{\theta}-\eta\nabla\mathcal{L}_{\mathrm{s},\mathbf{p}}(\bm{\theta})$ is real-analytic, its Jacobian $D\phi(\bm{\theta})=\mathbf{I}_{p}-\eta\nabla^{2}\mathcal{L}_{\mathrm{s},\mathbf{p}}(\bm{\theta})$ is real-analytic, and so is $\bm{\theta}\mapsto\det D\phi(\bm{\theta})$ (the determinant is a polynomial in the matrix entries).
Step 2: Witness and measure-zero critical set.
We rule out the degenerate case by a witness: at $\bm{\theta}_{\star}=\mathbf{0}_{p}$, our Hessian calculation (Lemma C.4) shows $\det D\phi(\bm{\theta}_{\star})>0$, hence $\det D\phi$ is not identically zero and its zero set $\mathcal{C}:=\{\det D\phi=0\}$ has Lebesgue measure zero by the real-analytic zero–set theorem (Theorem A.1; summarized in Theorem C.3).
Step 3: Local-to-global via countable chart covers.
On the complement $\mathbb{R}^{p}\setminus\mathcal{C}$, the Inverse Function Theorem (Theorem A.2) provides, for every $\bm{\theta}$, a neighborhood on which $\phi$ is a $C^{1}$ diffeomorphism. Although these neighborhoods form an a priori uncountable cover, the second countability of $\mathbb{R}^{p}$ (and of its subspaces) ensures a countable subcover of such charts (Proposition A.15, Lemma C.5). This countability is crucial because it lets us pass from local statements to a global measure statement via countable unions. With this cover in hand, the change-of-variables formula on each chart (Theorem C.4) implies that the image under the local inverse of any null set remains null; piecing the charts together and adding the null set $\mathcal{C}$ shows that preimages of Lebesgue-null sets under $\phi$ are null (Lemma C.6).
Step 4: Preservation of absolute continuity and conclusion.
Equivalently, $\phi$ pushes absolutely continuous laws to absolutely continuous laws (Theorem C.5); iterating across finitely many GD steps preserves absolute continuity (Corollary C.5.1). Finally, combining this preservation with the almost-sure pairwise distinctness of last-token representations over any finite input family (Corollary C.2.1) yields the main consequence we need for training: the last-token representation map remains injective almost surely after any finite GD horizon.
@@ C.2.1 Witness Construction
The goal of this subsubsection is to show that the GD Jacobian determinant is not identically zero by evaluating it at the all-zeros witness $\bm{\theta}_{\star}=\mathbf{0}_{p}$. The argument proceeds bottom-up: we first establish a “zero-gate” lemma that zeroes out most Hessian blocks when the gate matrix vanishes (Lemma C.1), derive the resulting spectrum (Lemma C.2, Lemma C.3), and then assemble these into the full Hessian at the witness (Lemma C.4).
@@
In particular, $Df(\mathbf{v}_{0})\big[(\mathbf{0}_{m},\;\cdot\;)\big]=\mathbf{0}_{n}$. The second-order chain rule for Fréchet derivatives (e.g. Magnus and Neudecker 2019, Thm. 18.4) yields:
@@
The determinant of a block triangular (in particular block diagonal) matrix equals the product of the determinants of its diagonal blocks (e.g. Horn and Johnson 2013, Cor. 0.8.5). Hence
@@
Lemma C.3 (Hessian of $\mathcal{L}$ w.r.t. $\mathbf{U},\bm{\beta}$ at $\bm{\theta}_{\star}=\mathbf{0}$ and its spectrum).
@@
Using $\mathrm{vec}(\mathbf{A}\mathbf{X}\mathbf{b})=(\mathbf{b}^{\top}\otimes\mathbf{A})\,\mathrm{vec}(\mathbf{X})$ (standard identity for vectorization, cf. Henderson and Searle (1981)), with $\mathbf{A}=\mathbf{I}_{n}$ and $\mathbf{b}=\bm{\beta}$,
@@
Combining, Lemmas C.2 and C.3, we get that
$\mathrm{spec}\big(\nabla^{2}\mathcal{L}(\bm{\theta}_{\star})\big)$ $=\mathrm{spec}\big(\nabla^{2}_{(\mathbf{u},\bm{\beta})}\mathcal{L}(\bm{\theta}_{\star})\big)\ \cup\ \{0\}^{\,d+p^{\prime}}$
@@
Then the critical set
@@
By Propositions B.3 and A.6 and the closure properties of real analyticity, $\mathcal{L}_{\mathrm{s},\mathbf{p}}$ is real-analytic; hence so are its gradient and Hessian. Therefore $\phi$ is real-analytic (Lewis, 2014, Thm. 1.1.15) and
@@
$\det D\phi(\bm{\theta}_{\star})=\left(1-\eta^{2}\|\mathbf{w}\|_{2}^{2}\right)^{d}>0.$
@@ C.2.2 Gradient Descent preserves absolute continuity
With the witness in hand and the critical set shown to be measure-zero (Theorem C.3), we now carry out Steps 3–4 of the roadmap: cover the non-critical region by countably many diffeomorphic charts (Lemma C.5), use the change-of-variables formula to show preimages of null sets remain null (Lemma C.6), and conclude that one GD step preserves absolute continuity (Theorem C.5). Iterating yields the finite-horizon corollary (Corollary C.5.1).
@@
1) $\bm{\mathcal{X}}$ is open: By Propositions B.3 and A.6 and the closure rules of real-analyticity, $\mathcal{L}_{\mathrm{s},\mathbf{p}}$ is $C^{2}$, hence $\phi$ is $C^{1}$. The map $\bm{\theta}\mapsto D\phi(\bm{\theta})$ is continuous, and the determinant is a continuous polynomial in the entries, so $g(\bm{\theta}):=\det D\phi(\bm{\theta})$ is continuous. Therefore $\mathcal{C}=g^{-1}(\{0\})$ is closed (Rudin, 1976, Thm. 4.8) and $\bm{\mathcal{X}}=\mathbb{R}^{p}\setminus\mathcal{C}$ is open.
@@
with stepsize $\eta\in(0,1)$. Then, gradient-descent preserves absolute continuity: for every absolutely continuous probability law $\mu$ on $\mathbb{R}^{p}$, its image under $\phi$ remains absolutely continuous:
@@
Appendix D Left-Invertibility Via SipIt
Goal.
We study when and how the hidden states of a causal decoder-only Transformer admit a left inverse: given the layer-$\ell$ representation at position $t$ and the true prefix $\pi=\mathrm{s}_{1:t-1}$, can we recover the next token $\mathrm{s}_{t}$?
Main idea.
Under mild randomness in the parameters and causal masking, the one-step last-token map that sends a candidate token $v$ to the layer-$\ell$ representation at position $t$ (conditioning on $\pi$) is almost-surely injective, and in fact has a positive separation margin. This yields a simple verifier: declare $v$ correct iff the observed hidden state lies in a small ball around $F(v;\pi,t)$.
Algorithmic consequence.
Because causality localizes the dependence to $(\pi,\mathrm{s}_{t})$, we can invert an entire sequence sequentially with a single pass over the vocabulary per position. We call this procedure SipIt (Sequential Inverse Prompt via ITerative updates), and we show exact (and robust) recovery holds almost surely, with worst-case time $\Theta(T|\mathcal{V}|)$.
@@ Appendix D Left-Invertibility Via SipIt
Every attention layer in every block is causal in the sense of Definitions B.6 and B.7. Consequently, for any $\mathrm{s}$ and any $t\in[T]$,
@@
SipIt is applied to models initialized with parameters drawn from an absolutely continuous distribution and trained via (mini-batch) gradient descent with step sizes in $(0,1)$, as described in Appendix C. Under these conditions, any network considered in the sequel is almost-surely injective (Theorem C.1).
@@ D.2 The Core Routines: Local Verifiers, Acceptance Regions, and Policies
Candidate enumeration.
Finally, we introduce the last conceptual block required to build the inversion algorithm: a policy algorithm that systematically enumerates candidate tokens so that the verifier is guaranteed to encounter the true one.
@@
Both Algorithms 2 and 3 satisfy Definition D.3. In practice we use the gradient-guided policy with standard gradient descent updates, as it tends to find the verified token with far fewer proposals: the next token is chosen by ranking $\mathcal{V}$ by the distance $\|\mathbf{E}_{v}-\mathbf{e}^{(j)}\|_{2}$ to the updated proxy $\mathbf{e}^{(j)}$. This preserves the same worst-case guarantees (single pass over $\mathcal{V}$) while improving empirical efficiency.
D.3 Global Inversion via SipIt
@@
We are now ready to introduce our inversion algorithm: SipIt (Algorithm 1). The algorithms applies to decoder-only transformers with causal self-attention (Assumption D.1), and assumes injectivity, which occurs with almost-surely (Assumption D.2). We assume access to the layer-$\ell$ hidden states per position $\left\{\widehat{\mathbf{h}}_{t}\right\}_{t=1}^{T}$ and to the parameters needed to evaluate the local verifier from Definition D.2 for arbitrary $(t,\pi,j)$, as well as the gradient (when needed), namely to the model up to layer $\ell$. A policy algorithm is fixed (e.g., Algorithm 3).
@@
Fix $t$ and a true prefix $\pi=\langle\mathrm{s}_{1},\ldots,\mathrm{s}_{t-1}\rangle$. Under Assumption D.1, the observed hidden state at step $t$ satisfies $\mathbf{h}_{t}(\mathrm{s})=F(\mathrm{s}_{t}\,;\,\pi,t)$ (Lemma D.2). In addition, under Assumption D.2, $F$ is injective and has positive margin $\Delta_{\pi,t}>0$ almost-surely (Theorems D.1 and D.1). Consequently, for the local verifier of Definition D.2, the following hold with probability one:
@@
Work on the probability-one event of Proposition D.2 (under Assumptions D.1 and D.2 with the stated $\varepsilon$), on which the local verifier at step $t$ uniquely accepts the true token $\mathrm{s}_{t}$. Equivalently,
@@
Theorem D.2 (Correctness of SipIt (noiseless & robust)).
For each $t\in\{1,\ldots,T\}$ let $\pi_{t}=\langle\mathrm{s}_{1},\ldots,\mathrm{s}_{t-1}\rangle$ and let $\Delta_{\pi_{t},t}>0$ be the margin of the one-step map $F(\cdot;\pi_{t},t)$ from Lemma D.1. Under Assumptions D.1 and D.2, run SipIt (Algorithm 1) with a tolerance $\varepsilon\geq 0$ and observations
@@
By Assumptions D.2 and D.1, and Lemma D.1, there is a probability-one event on which, for all $t$, $F(\cdot;\pi_{t},t)$ is injective with strictly positive margin $\Delta_{\pi_{t},t}$. Intersecting across finitely many $t$ preserves probability 1. Work on this event.
By Assumptions D.1 and D.2, $\mathbf{h}_{t}(\mathrm{s})=F(\mathrm{s}_{t};\pi_{t},t)$. Since $\|\mathbf{e}_{t}\|_{2}\leq\varepsilon$,
@@
When $\varepsilon=0$, equation 36 also holds by Proposition D.1(1). We now analyze SipIt and proceed by induction on $t$.
@@
Termination and complexity.
Having established correctness, we record the worst-case iteration count and discuss its relation to wall-clock time.
@@
Run SipIt (Algorithm 1) on a length-$T$ sequence with any policy that enumerates $\mathcal{V}$ without replacement. Then the algorithm halts after a finite number of iterations. Moreover, in the worst case the inner for-loop over $j$ executes at most $|\mathcal{V}|$ iterations at each position $t$, so the total number of verifier tests across the entire run is at most $T\,|\mathcal{V}|$. In particular, the number of loop iterations grows linearly with $T\cdot|\mathcal{V}|$.
@@
Appendix E Implementation Details and Additional Experiments
This section collects implementation details (§E.1), additional ablations on collision experiments and SipItE.2), controlled experiments with identical next-token predictions (§E.3), qualitative inspection of the closest hidden-state pairs (§E.4), and connections to anisotropy and intrinsic dimension (§E.5).
E.1 Implementation Details
What is a collision in practice.
In the theoretical parts of the paper we use “collision” in the usual functional sense: two distinct prompts $\mathrm{s}\neq\mathrm{s}^{\prime}$ such that their last-token representations coincide exactly,
$\mathbf{r}(\mathrm{s}\,;\,\bm{\theta}_{T})=\mathbf{r}(\mathrm{s}^{\prime}\,;\,\bm{\theta}_{T}).$
This is the event whose probability is controlled in theorems 2.2 and 2.3 and in Appendix C, and all proofs are carried out at the level of exact equality (no numerical threshold is required).
In the experiments, however, representations are stored in floating-point format, so exact equality of $\mathbf{r}(\mathrm{s}\,;\,\bm{\theta}_{T})$ and $\mathbf{r}(\mathrm{s}^{\prime}\,;\,\bm{\theta}_{T})$ may not be a meaningful or robust criterion. We therefore adopt a numerical proxy: given two prompts $\mathrm{s},\mathrm{s}^{\prime}$ and their embeddings $\mathbf{r}(\mathrm{s}\,;\,\bm{\theta}_{T}),\mathbf{r}(\mathrm{s}^{\prime}\,;\,\bm{\theta}_{T})\in\mathbb{R}^{d}$, we declare a practical collision only if
$\texttt{torch.allclose}\bigl(\mathbf{r}(\mathrm{s}\,;\,\bm{\theta}_{T}),\mathbf{r}(\mathrm{s}^{\prime}\,;\,\bm{\theta}_{T})\bigr)=\texttt{True},$
i.e., every coordinate falls within PyTorch’s prescribed tolerances, namely $10^{-5}$ and $10^{-8}$ for relevant and absolute tolerance respectively. Across all of the billions to trillions of empirical checks, every pair of distinct prompts $\mathrm{s}\neq\mathrm{s}^{\prime}$ failed this criterion: torch.allclose always returned False, and the observed $\ell_{2}$ distances were consistently bounded away from zero. Thus, at the resolution of our numerical precision, we did not observe any collisions in practice.
SipIt implementation.
We implement SipIt exactly as in Alg. 1 with the gradient-guided policy. To stabilize the continuous proxy used for ranking, we apply gradient clipping (capping the gradient norm at 1) and we periodically project it back to the nearest token embedding every K=50K\!=\!50 candidate proposals:
without taking gradients through this projection. These heuristics affect efficiency only; the verifier and all correctness guarantees remain unchanged.
HardPrompts implementation.
The original HardPrompts method Wen et al. (2023) targets multimodal vision-language models and optimizes prompts via a CLIP-based similarity objective. In our text-only setting we lack the vision branch and CLIP loss, so we adapt Algorithm 1 of Wen et al. (2023) to language models by replacing the objective with the same $\ell_{2}$ loss used in SipIt’s gradient calculation, and setting the optimization steps $T=\tfrac{1}{4}\text{\# tokens}\cdot|\mathcal{V}|$. All other details (step sizes, stopping rules) mirror our SipIt setup to ensure a fair comparison.